5. Solve the following equations in polar form and plot the roots in the complex plane:
(a) x6 = 1 (b) 24 = -1 (c) 24 = 1+iV3
Answer #1

4 3 = g(25T+9) , this j(2nT) 13 = y gi (me), n- 0) 2,3. 4 ) - ) ④ = = | 3 ) (17-122.4 - 0 3 | C e racious of circle = 131 il© 34= Hiss Efiam hic in = 2(+iva) = 2 (Costa Sinta) → 34= 2 eise 2e (2nT+5) 3= 240 e 1 (2+ ), n=0,1,2,3 =244 e il ), n = 0,1,